# Lattice QCD Benchmark of Proton Helicity and Flavor-Dependent Unpolarized TMDPDFs at Physical Quark Masses

### v1

@@ Abstract
We present the first lattice QCD calculations of the isovector helicity transverse momentum-dependent parton distribution function (TMDPDF) and the flavor-dependent unpolarized TMDPDFs for up and down quarks in the proton. Our computations utilize domain-wall fermion discretization with physical quark masses. Employing Coulomb-gauge-fixed bilocal quark correlation functions within the large-momentum effective theory framework, we access nonperturbative transverse quark separations $b_{T}$ up to approximately 1 fm, corresponding to low transverse momentum scales. We present renormalization and factorization scale-independent ratios of these TMDPDFs as functions of $b_{T}$ and longitudinal momentum fraction $x$, allowing direct comparisons with phenomenological parameterizations from global experimental fits. Our results demonstrate remarkably similar $b_{T}$ dependence between helicity and unpolarized TMDPDFs at moderate $x$, in agreement with phenomenology. In contrast to certain phenomenological models, we observe mild flavor dependence in the $b_{T}$ distributions at moderate values of $x$.
I Introduction
Protons, as the stable constituents of atomic nuclei, form the basis of visible matter. Understanding their internal structure and dynamics from first principles of Quantum Chromodynamics (QCD), the theory of strong interactions, is a central goal of modern theoretical and experimental physics. This requires unraveling how the proton’s spin emerges from its quark and gluon constituents. While collinear parton distribution functions (PDFs) provide information on longitudinal momentum structure, a complete picture of proton spin demands access to its full three-dimensional momentum structure. Transverse momentum dependent PDFs (TMDPDFs) generalize PDFs by incorporating dependence on both the longitudinal momentum fraction $x$ and transverse momentum $k_{T}$. This makes TMDPDFs central to the science goals of the upcoming Electron-Ion Collider (EIC) Boer et al. (2011); Accardi et al. (2016); Abdul Khalek et al. (2022); Abir et al. (2023), which aims to probe the proton’s spin structure with unprecedented precision.
@@ I Introduction
Phenomenological parameterizations of TMDPDFs from global fits to semi-inclusive deep inelastic scattering and Drell-Yan data have made substantial progress in characterizing their nonperturbative structure Davies et al. (1984); Ladinsky and Yuan (1994); Landry et al. (2003); Konychev and Nadolsky (2006); Sun et al. (2018); D’Alesio et al. (2014); Bacchetta et al. (2017); Scimemi and Vladimirov (2018); Bertone et al. (2019); Scimemi and Vladimirov (2020); Bacchetta et al. (2020); Hautmann et al. (2020); Bury et al. (2022); Bacchetta et al. (2022); Isaacson et al. (2023); Aslan et al. (2024); Vladimirov (2019); Cerutti et al. (2023); Barry et al. (2023) including recent studies of flavor dependence Moos et al. (2024); Yang et al. (2024); Bacchetta et al. (2024a). These approaches are limited by sparse data and rely on model-dependent assumptions for nonperturbative inputs at low $k_{T}$ , where there are still tensions among different groups. Global fits of the proton helicity TMDPDF have only become available very recently Yang et al. (2024); Bacchetta et al. (2024b), but the results from the two groups show noticeable inconsistencies.
@@
Lattice QCD provides a complementary, first-principles approach by enabling direct calculation of spin- and flavor-resolved parton structure in the nonperturbative regime. While light-cone TMDPDFs are defined in Minkowski spacetime and inaccessible to Euclidean lattice QCD, they can be approached within the large-momentum effective theory framework Ji (2013, 2014); Ji et al. (2021a) via quasi-TMDPDFs Ji et al. (2015, 2019); Ebert et al. (2019a, b); Ji et al. (2020a, b); Ebert et al. (2020); Vladimirov and Schäfer (2020); Ji et al. (2021b); Ji and Liu (2022); Ebert et al. (2022); Schindler et al. (2022); Zhu et al. (2023); Rodini and Vladimirov (2023). At large nucleon momentum, quasi-TMDPDFs are matched to their light-cone counterparts through perturbative factorization Ebert et al. (2019b); Ji et al. (2020a, b); Ebert et al. (2022), enabling access to observables such as the CS kernel Shanahan et al. (2020, 2021); Schlemmer et al. (2021); Shu et al. (2023); Zhang et al. (2020); Li et al. (2022); Chu et al. (2022, 2023a); Avkhadiev et al. (2023, 2024); Bollweg et al. (2024, 2025), soft functions Zhang et al. (2020); Li et al. (2022); Chu et al. (2023a); Bollweg et al. (2025), and nucleon and pion TMDPDFs He et al. (2022); Chu et al. (2023b); Walter et al. (2024); Ma et al. (2025); Bollweg et al. (2025).
@@
Traditional lattice calculations of quasi-TMDPDFs use staple-shaped Wilson lines to preserve gauge invariance. These introduce power-divergent self-energies and substantial statistical noise, especially at large $b_{T}$, limiting precision. The recently developed Coulomb-gauge-fixed (CG) quasi-TMDPDF approach Gao et al. (2024a); Zhao (2023) addresses these issues by eliminating Wilson lines entirely. This reduces noise and simplifies renormalization. The CG method has enabled improved CS kernel extractions Bollweg et al. (2024); Mukherjee et al. (2024) and the first lattice determination of the pion TMDPDF with the soft function Bollweg et al. (2025), demonstrating its potential for advancing lattice TMD studies.
@@
In this work, we present the first lattice QCD calculation of the isovector helicity and flavor-dependent unpolarized TMDPDFs in the proton taking advantage of the CG quasi-TMDPDF framework. The calculation employs physical quark masses and chiral-symmetry-preserving domain-wall fermion discretization, reaching the nonperturbative region up to $b_{T}\approx 1~{}\mathrm{fm}$. By forming renormalization- and factorization-scale-independent ratios, we enable direct comparisons with phenomenological fits without requiring separate determinations of soft functions. Our results show a remarkable similarity between the isovector helicity and unpolarized TMDPDFs at moderate $x$, and reveal only mild flavor dependence in the unpolarized TMDPDFs—providing the first lattice benchmark in these channels. These findings offer valuable nonperturbative constraints for global analyses, help resolve tensions among existing fits, and establish a solid foundation for future lattice studies of the proton’s spin structure.
II Theoretical framework
@@
The bare quasi-TMD beam functions are defined as matrix elements of bilocal quark bilinears with spatial separations in both the transverse ($b_{T}$) and longitudinal ($b_{z}$) directions, sandwiched between boosted hadron states. For CG quasi-TMDs, the definition is
@@
with $\mathcal{O}_{\Gamma}(b_{T},b_{z})=\bar{\psi}(b_{T},\frac{b_{z}}{2})\Gamma\psi(0,-\frac{b_{z}}{2})|_{\nabla\cdot\textbf{A}=0}$. Here, $\psi$ is the quark field, $\nabla\cdot\textbf{A}=0$ is the CG condition, $P_{z}$ is the hadron momentum, $\lambda$ the hadron polarization, and $\Gamma$ specifies the TMDPDF structure. This contrasts with the gauge-invariant (GI) quasi-TMD beam function, where non-local quark fields are connected by staple-shaped Wilson lines to preserve gauge invariance (see, e.g., Ref. Ji et al. (2020b); Ebert et al. (2022)).
After appropriate renormalization of $\tilde{h}^{B}$ (explained later in the text) and Fourier transforming the renormalized operator $\tilde{h}^{R}$ with respect to $b_{z}$ one obtains the quasi-TMD beam function $\tilde{f}(x,b_{T},P_{z},\mu)$, which is related to the light-cone TMDPDF $f(x,b_{T},\zeta,\mu)$ via the factorization Ebert et al. (2022),
$\begin{split}&\frac{\tilde{f}(x,b_{T},P_{z},\mu)}{\sqrt{S_{r}(b_{T},\mu)}}=H(x,P_{z},\mu)e^{\frac{1}{2}\ln\frac{(2xP_{z})^{2}}{\zeta}\gamma^{\overline{\rm MS}}(b_{T},\mu)}f(x,b_{T},\zeta,\mu)+\mathcal{O}\left(\frac{\Lambda_{\rm QCD}^{2}}{(xP_{z})^{2}},\frac{1}{(b_{T}(xP_{z}))^{2}}\right),\end{split}$ (2)
with power corrections suppressed by $1/(xP_{z})$. Here, $\mu$ and $\zeta$ are the factorization and rapidity scales. The evolution is governed by perturbative anomalous dimensions and the CS kernel $\gamma^{\overline{\rm MS}}(b_{T},\mu)$. $H(x,P_{z},\mu)$ is the hard kernel, and $S_{r}(b_{T},\mu)$ is the intrinsic soft factor Ji et al. (2020a).
This factorization is independent of quark flavor and spin structure, and involves no gluon mixing Ebert et al. (2022). These features enable the construction of renormalization and factorization scale-independent renormalization-group-invariant (RGI) ratios. RGI ratios cancel the intrinsic soft factor and correspond to light-cone TMDPDF ratios to all orders in perturbation theory, up to power corrections.
For example, the ratio of isovector helicity TMDPDF $g_{1L}^{\Delta u_{+}-\Delta d_{+}}$ to isovector unpolarized TMDPDF $f_{1}^{u_{v}-d_{v}}$ of proton can be expressed as
$\begin{split}R^{u-d}_{g_{1L}/f_{1}}(x,b_{T})\cdot g_{A}&\equiv\frac{g_{1L}^{\Delta u_{+}-\Delta d_{+}}(x,b_{T},\zeta,\mu)}{f_{1}^{u_{v}-d_{v}}(x,b_{T},\zeta,\mu)}=\frac{\tilde{g}_{1L}^{\Delta u_{+}-\Delta d_{+}}(x,b_{T},P_{z},\mu)}{\tilde{f}_{1}^{u_{v}-d_{v}}(x,b_{T},P_{z},\mu)},\end{split}$ (3)
@@ III Lattice QCD calculations
The quasi-TMD beam function matrix elements are extracted from proton three-point correlation functions,
@@
$\begin{split}&C^{\mathrm{3pt}}_{\mathcal{P},\Gamma}(t_{s},\tau;b_{T},b_{z},P_{z})=\sum_{\vec{y},\vec{x}}e^{-iP_{z}(y_{z}-x_{z})}\mathcal{P}^{\mathrm{3pt}}_{\alpha\beta}\langle N_{\alpha}(\vec{y},t_{s}+t)\mathcal{O}_{\Gamma}(b_{T},b_{z},\tau+t)\overline{N}_{\beta}(\vec{x},t)\rangle.\end{split}$ (5)
@@
We use the proton interpolating field $N^{s}_{\alpha}(\vec{x},t)=\varepsilon_{abc}u^{s}_{a\alpha}(\vec{x},t)\left[u^{s}_{b}(\vec{x},t)^{T}(\gamma_{t}\gamma_{y})\gamma_{5}d^{s}_{c}(\vec{x},t)\right]$, where the superscript $s$ denotes smeared quark fields. The source position of $N$ is denoted by $(\vec{x},t)$, $t_{s}$ is the source-sink separation, $\tau$ is the operator insertion time, and $\mathcal{P}^{\mathrm{3pt}}_{\alpha\beta}$ selects the nucleon spin polarization. For the helicity quasi-TMDPDF, we use $\Gamma=i\gamma_{5}\gamma_{z}$ and $\mathcal{P}^{\mathrm{3pt}}_{\alpha\beta}=\frac{1}{2}(1+\gamma_{t})i\gamma_{5}\gamma_{z}$; for the unpolarized quasi-TMDPDF, we use $\Gamma=\gamma_{t}$ and $\mathcal{P}^{\mathrm{3pt}}_{\alpha\beta}=\frac{1}{2}(1+\gamma_{t})$.
@@
Calculations are performed using 81 gauge-field configurations of the 2+1-flavor domain-wall fermion ensemble 64I Blum et al. (2023), with 5-dimensional volume $N_{s}^{3}\times N_{t}\times L_{5}=64^{3}\times 128\times 12$, physical up, down, and strange quark masses, and lattice spacing $a=0.0836$ fm.
@@
To improve ground-state overlap and reduce statistical noise, we employ boosted Gaussian smearing Bali et al. (2016) of the quark fields in $N$. Gaussian smearing is implemented with radius $r_{G}=0.75$ fm and quark boost parameter $j_{z}=0$ and 3 Gao et al. (2021, 2020), corresponding to nucleon momentum $P_{z}=2\pi n_{z}/(N_{s}a)$ with $n_{z}=0$ and 7, yielding $P_{z}^{\mathrm{max}}=1.62$ GeV.
@@
Quark propagators are computed from CG-fixed gauge configurations using a deflated solver with 2000 eigenvectors. For each gauge configuration, we perform multiple measurements of $C^{\mathrm{3pt}}_{\mathcal{P},\Gamma}$ using the all-mode averaging technique Shintani et al. (2015). For $n_{z}=0$, we compute 1 exact solve (solver tolerance $10^{-8}$) and 4 sloppy solves (solver tolerance $10^{-4}$) at $t_{s}/a=\{6,8,10\}$. For $n_{z}=7$, we compute 1 exact and {16, 32, 64, 128} sloppy solves at $t_{s}/a=\{4,6,8,10\}$, respectively.
@@
We compute isovector ($u-d$) contributions for both helicity and unpolarized quasi-TMDPDFs. To study flavor dependence, the isoscalar ($u+d$) channel is also calculated for the unpolarized case, neglecting quark-line disconnected contributions which is usually mild but need to be studied in future work.
For completeness and comparison, gauge-invariant quasi-TMDPDFs are also computed. Details of these calculations are provided in Appendix C.
@@
Figure 1: Ratios $r(t_{s},\tau;b_{T},b_{z},P_{z}=1.62~{}\mathrm{GeV})$ for unpolarized (left) and helicity (right) CG quasi-TMD beam function matrix elements at $b_{T}=8a$ and $b_{z}=0$. Dashed lines show results from two-state fits; bands indicate the extracted values from $\mathrm{Fit}_{\mathrm{st1}}$, $\mathrm{Fit}_{\mathrm{st2}}$, and SumFit.
IV Results
@@
We extract the bare matrix elements of the CG quasi-TMD beam functions by analyzing the ratio of three-point to two-point correlators,
@@
where $C^{\mathrm{2pt}}$ denotes the two-point correlator of the unpolarized proton. In the limit $t_{s}>\tau\rightarrow\infty$, excited-state contributions are exponentially suppressed, and the ratio isolates the ground-state matrix elements $\tilde{h}^{B}(b_{T},b_{z},P_{z},a)$. For finite $t_{s}$ and $\tau$, we apply three extraction strategies: one-state/plateau fit ($\mathrm{Fit}_{\mathrm{st1}}$), two-state fit ($\mathrm{Fit}_{\mathrm{st2}}$), and summation fit (SumFit).
@@
In $\mathrm{Fit}_{\mathrm{st1}}$, we fit a constant to the stable region of $r$ for $t_{s}=6a,8a,10a$, excluding up to $\tau=3a$ at both ends of each source-sink separation. For $\mathrm{Fit}_{\mathrm{st2}}$, we use $\tau\in[2a,t_{s}-2a]$ and truncate the spectral sum in both $C^{\mathrm{2pt}}$ and $C^{\mathrm{3pt}}$ to include only the ground and first excited states. We extract two energy levels and corresponding overlap amplitudes from fits to $C^{\mathrm{2pt}}$, and then use these parameters in the spectral decomposition of $r$ to determine the matrix elements. For SumFit, we use $r$ at $t_{s}=4a,6a,8a,10a$, summing over $\tau\in[2a,t_{s}-2a]$, and extract the matrix element from the slope of a linear fit in $t_{s}$.
@@
Figure 1 shows representative results at $P_{z}=1.62~{}\mathrm{GeV}$, $b_{T}=8a$, and $b_{z}=0$, for unpolarized (left) and helicity (right) matrix elements. Consistency across the methods indicates that excited-state effects are under control. For the final analysis, we conservatively average over all three methods on a bootstrap sample-by-sample basis. Additional details and examples are provided in Appendix A.
Figure 2: Real parts of renormalized matrix elements for CG quasi-TMD beam functions at nucleon momentum $P_{z}=1.62$ GeV, and for differnt $b_{T}$, shown as functions of $b_{z}$. From left to right: isoscalar unpolarized, isovector unpolarized, and isovector helicity channels. Bands indicate exponential extrapolations from fits with $b_{z}^{L}=16a$.

### v3

# Lattice QCD Benchmark of Proton Helicity and Flavor-Dependent Unpolarized Transverse Momentum-Dependent Parton Distribution Functions at Physical Quark Masses

@@ Abstract
We present the first lattice QCD calculations of the isovector helicity transverse momentum-dependent parton distribution function (TMDPDF) and the flavor-dependent unpolarized TMDPDFs for up and down quarks in the proton. Our computations utilize domain-wall fermion discretization with physical quark masses. Employing Coulomb-gauge-fixed quark correlation functions within the large-momentum effective theory framework, we access nonperturbative transverse quark separations $b_{T}$ up to approximately 1 fm, corresponding to transverse momenta as low as 200 MeV. Based on the quasi-TMD factorization theorem, we construct renormalization-group–invariant ratios that are equal to the corresponding light-cone TMDPDF ratios. At moderate $x$, our results reveal that the isovector helicity and unpolarized TMDPDFs exhibit nearly identical transverse structure up to a normalization factor, and the unpolarized distributions display only mild flavor dependence. These findings not only support key trends observed in recent global analyses but also provide robust, nonperturbative constraints that can distinguish between different parametrizations. This work establishes a first-principles benchmark for TMDPDFs, offering valuable input for ongoing and future experimental efforts to map the proton’s three-dimensional structure.
Introduction. Protons, as the stable constituents of atomic nuclei, form the basis of visible matter. Understanding their internal structure and dynamics from first principles of Quantum Chromodynamics (QCD), the theory of strong interactions, is a central goal of modern theoretical and experimental physics. This requires unraveling how the proton’s spin emerges from its quark and gluon constituents. While collinear parton distribution functions (PDFs) provide information on longitudinal momentum structure, a complete picture of proton spin demands access to its full three-dimensional momentum structure. Transverse momentum dependent PDFs (TMDPDFs) generalize PDFs by incorporating dependence on both the longitudinal momentum fraction $x$ and transverse momentum $k_{T}$. This makes TMDPDFs central to the scientific objectives of experiments at facilities such as Jefferson Lab [1, 2], the Large Hadron Collider (LHC)[3, 4], and the upcoming Electron-Ion Collider (EIC)[5, 6, 7, 8], which will explore the proton’s spin structure with unprecedented precision.
@@
Phenomenological parametrizations of the unpolarized TMDPDF from global fits to semi-inclusive deep inelastic scattering and Drell-Yan data have made substantial progress in the past four decades [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], including recent studies of its flavor dependence [30, 31]. This approach is limited by sparse data in the kinematic region sensitive to nonperturbative physics and rely on model-dependent assumptions at low $k_{T}$ , where there are still tensions among different groups. In contrast, global fits of the proton helicity TMDPDF have only become available very recently [32, 33], but the results from the two groups show noticeable inconsistencies.
@@
Lattice QCD is a first-principles approach that can calculate the spin- and flavor-resolved hadron structure in the nonperturbative regime. Although TMDPDFs are defined from time-dependent light-cone correlations and inaccessible on a Euclidean lattice, they can be computed within the large-momentum effective theory framework [34, 35, 36] via the quasi-TMDPDFs [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. At large hadron momentum, the quasi-TMDPDFs are matched to their light-cone counterparts through an effective theory expansion [40, 41, 42, 47], enabling access to observables such as the Collins-Soper (CS) kernel for TMD evolution [51, 52, 53, 54, 55, 56, 57, 58, 59, 60], soft function [55, 56, 58], and nucleon and pion TMDPDFs [61, 62, 63, 64].
@@
Traditional lattice calculations of quasi-TMDPDFs use nonlocal quark bilinears with staple-shaped Wilson lines, which introduce linearly power divergence and substantial statistical noise, especially at large $b_{T}$ (or equivalently small $k_{T}$), thus limiting the precision. The recently developed Coulomb-gauge (CG) quasi-TMDPDF approach [65, 66] addresses these issues by eliminating Wilson lines entirely, which not only reduces the statistical noise but also simplifies the renormalization. The CG method has enabled improved CS kernel calculations [67, 68] and the first lattice determination of the pion TMDPDF with the soft function [69], demonstrating its potential for advancing TMD studies.
@@
In this Letter, we present the first lattice QCD calculation of the isovector helicity and flavor-dependent unpolarized TMDPDFs in the proton, taking advantage of the CG method. The calculation employs chiral-symmetry-preserving domain-wall fermion discretization and physical quark masses, and reaches the nonperturbative region up to $b_{T}\approx 1~\mathrm{fm}$ which corresponds to $k_{T}\sim 200$ MeV. By forming renormalization- and factorization-scale-independent ratios, we enable direct comparisons with phenomenological fits without requiring a separate determination of the soft function. Our results show a remarkable similarity between the isovector helicity and unpolarized TMDPDFs at moderate $x$, and reveal only mild flavor dependence in the unpolarized TMDPDFs—providing the first lattice benchmark in these channels. These findings offer valuable nonperturbative constraints for global analyses and establish a solid foundation for future lattice studies of the proton’s 3D spin structure.
@@
Theoretical framework. The bare CG quasi-TMD beam functions are defined as matrix elements of nonlocal quark bilinears with spatial separations in both the transverse ($b_{T}$) and longitudinal ($b_{z}$) directions, sandwiched between boosted hadron states,
@@
with $\mathcal{O}_{\Gamma}(b_{T},b_{z})=\bar{\psi}(b_{T},\frac{b_{z}}{2})\Gamma\psi(0,-\frac{b_{z}}{2})|_{\nabla\cdot\textbf{A}=0}$. Here, $\psi$ is the quark field, $\nabla\cdot\textbf{A}=0$ is the CG condition, $P_{z}$ is the hadron momentum, $\lambda$ the hadron polarization, and $\Gamma$ specifies the TMDPDF spin structure. This contrasts with the gauge-invariant (GI) quasi-TMD beam function, where non-local quark fields are connected by staple-shaped Wilson lines to preserve gauge invariance (see, e.g., Ref. [37]).
After appropriate renormalization of $\tilde{h}^{B}$—as explained below—and Fourier transform (FT) of the renormalized operator $\tilde{h}^{R}$ with respect to $b_{z}$, one obtains the quasi-TMD beam function $\tilde{f}(x,b_{T},P_{z},\mu)$, which is related to the light-cone TMDPDF $f(x,b_{T},\zeta,\mu)$ via the factorization formula [40, 41, 42, 47, 66],
$\begin{split}&\frac{\tilde{f}(x,b_{T},P_{z},\mu)}{\sqrt{S_{r}(b_{T},\mu)}}=H(xP_{z},\mu)\exp\left[{\frac{\gamma^{\overline{\rm MS}}(b_{T},\mu)}{2}\ln\frac{(2xP_{z})^{2}}{\zeta}}\right]\\ &\times f(x,b_{T},\zeta,\mu)+\mathcal{O}\left(\frac{\Lambda_{\rm QCD}}{xP_{z}},\frac{1}{b_{T}(xP_{z})}\right),\end{split}$ (2)
with power corrections suppressed by $1/(xP_{z})$ [70], so that the factorization is valid in the moderate-$x$ region but is expected to break down near the endpoints. Here, $\mu$ and $\zeta$ are the factorization and rapidity scales. The evolution in $\zeta$ is governed by the CS kernel $\gamma^{\overline{\rm MS}}(b_{T},\mu)$ [71, 3]. $H(xP_{z},\mu)$ is a hard kernel obtained at one-loop [66], and $S_{r}(b_{T},\mu)$ is the intrinsic soft factor [41].
This factorization is independent of quark flavor and spin and does not involve gluon mixing [47]. These features enable the construction of renormalization-group–invariant (RGI) ratios from quasi-TMD beam functions that are free from scheme and scale dependence. In particular, the cancellation of the soft factor, hard kernel, and scale evolution terms makes the RGI ratios identical to their light-cone counterparts to all orders in perturbation theory, up to power-suppressed corrections.
For example, the ratio of proton isovector helicity TMDPDF $g_{1L}^{\Delta u_{+}-\Delta d_{+}}$ to isovector unpolarized TMDPDF $f_{1}^{u_{v}-d_{v}}$ can be expressed as
$\begin{split}R^{u-d}_{g_{1L}/f_{1}}(x,b_{T})\cdot g_{A}&\equiv\frac{g_{1L}^{\Delta u_{+}-\Delta d_{+}}(x,b_{T},\zeta,\mu)}{f_{1}^{u_{v}-d_{v}}(x,b_{T},\zeta,\mu)}\\ &=\frac{\tilde{g}_{1L}^{\Delta u_{+}-\Delta d_{+}}(x,b_{T},P_{z},\mu)}{\tilde{f}_{1}^{u_{v}-d_{v}}(x,b_{T},P_{z},\mu)},\end{split}$ (3)
@@
Lattice QCD calculations. The quasi-TMD beam function matrix elements are extracted from proton three-point correlation functions,
@@
$\begin{split}&C^{\mathrm{3pt}}_{\mathcal{P},\Gamma}(t_{s},\tau;b_{T},b_{z},P_{z})=\sum_{\vec{y},\vec{x}}e^{-iP_{z}(y_{z}-x_{z})}\mathcal{P}^{\mathrm{3pt}}_{\alpha\beta}\\ &\times\langle N_{\alpha}(\vec{y},t_{s}+t)\mathcal{O}_{\Gamma}(b_{T},b_{z},\tau+t)\overline{N}_{\beta}(\vec{x},t)\rangle.\end{split}$ (5)
@@
We use the proton interpolating field $N^{s}_{\alpha}(\vec{x},t)=\varepsilon_{abc}u^{s}_{a\alpha}(\vec{x},t)\left[u^{s}_{b}(\vec{x},t)^{T}(\gamma_{t}\gamma_{y})\gamma_{5}d^{s}_{c}(\vec{x},t)\right]$, where the superscript $s$ denotes smeared quark fields. The source position of $N$ is denoted by $(\vec{x},t)$, $t_{s}$ is the source-sink separation, $\tau$ is the operator insertion time, and $\mathcal{P}^{\mathrm{3pt}}_{\alpha\beta}$ selects the nucleon spin polarization. For the helicity case, we use $\Gamma=i\gamma_{5}\gamma_{z}$ and $\mathcal{P}^{\mathrm{3pt}}_{\alpha\beta}=\frac{1}{2}(1+\gamma_{t})i\gamma_{5}\gamma_{z}$; for the unpolarized case, we use $\Gamma=\gamma_{t}$ and $\mathcal{P}^{\mathrm{3pt}}_{\alpha\beta}=\frac{1}{2}(1+\gamma_{t})$.
@@
The lattice calculations are performed using 81 gauge-field configurations of the 2+1-flavor domain-wall fermion ensemble 64I [72], with 5-dimensional volume $N_{s}^{3}\times N_{t}\times L_{5}=64^{3}\times 128\times 12$, physical up, down, and strange quark masses, and spacing $a=0.0836$ fm.
@@
To improve ground-state overlap and reduce statistical noise, we employ boosted Gaussian smearing [73] of the quark fields in the interpolation operator. Gaussian smearing is implemented with radius $r_{G}=0.75$ fm and quark boost parameter $j_{z}=0$ and 3 [74, 75], corresponding to nucleon momentum $P_{z}=2\pi n_{z}/(N_{s}a)$ with $n_{z}=0$ and 7, yielding $P_{z}^{\mathrm{max}}=1.62$ GeV.
@@
Quark propagators are computed from CG-fixed gauge configurations using a deflated solver with 2000 eigenvectors. For each gauge configuration, we perform multiple measurements of $C^{\mathrm{3pt}}_{\mathcal{P},\Gamma}$ using the all-mode averaging technique [76]. For $n_{z}=0$, we compute one exact solve [77, 78] (solver tolerance $10^{-8}$) and 4 sloppy solves (solver tolerance $10^{-4}$) at $t_{s}/a=\{6,8,10\}$. For $n_{z}=7$, we compute 1 exact and {16, 32, 64, 128} sloppy solves at $t_{s}/a=\{4,6,8,10\}$, respectively.
@@
We compute isovector (udu-d) contributions for both helicity and unpolarized quasi-TMD beam functions. To study flavor dependence, the isoscalar ($u+d$) channel is also calculated for the unpolarized case, neglecting quark-line disconnected contributions. These contributions are expected to be small [79, 80, 81], compared to the other systematic uncertainties. A dedicated investigation is needed in future work to quantify them.
@@
Figure 1: Ratios $r(t_{s},\tau;b_{T},b_{z},P_{z}=1.62~\mathrm{GeV})$ for unpolarized (top) and helicity (bottom) CG quasi-TMD beam function matrix elements at $b_{T}=8a$ and $b_{z}=0$. Dashed lines show results from two-state fits; bands indicate the extracted values from $\mathrm{Fit}_{\mathrm{st1}}$, $\mathrm{Fit}_{\mathrm{st2}}$, and SumFit.
@@
Results. We extract the bare matrix elements of the CG quasi-TMD beam functions by analyzing the ratio of three-point to two-point correlators,
@@
where $C^{\mathrm{2pt}}$ denotes the two-point correlator of the unpolarized proton. In the limit $\tau,t_{s}-\tau\rightarrow\infty$, the excited-state contributions are exponentially suppressed, and the ratio isolates the ground-state matrix elements $\tilde{h}^{B}(b_{T},b_{z},P_{z},a)$. For finite $t_{s}$ and $\tau$, we apply three extraction strategies: one-state/plateau fit ($\mathrm{Fit}_{\mathrm{st1}}$), two-state fit ($\mathrm{Fit}_{\mathrm{st2}}$), and summation fit (SumFit).
@@
In $\mathrm{Fit}_{\mathrm{st1}}$, we fit a constant to the stable region of $r$ for $t_{s}=6a,8a,10a$, excluding up to $\tau=3a$ at both ends of each source-sink separation. For $\mathrm{Fit}_{\mathrm{st2}}$, we use $\tau\in[2a,t_{s}-2a]$ for joint fit of $t_{s}\in[4a,10a]$ and truncate the spectral sum in both $C^{\mathrm{2pt}}$ and $C^{\mathrm{3pt}}$ to include only the ground and first excited states. We extract the two energy levels and corresponding overlap amplitudes from fits to $C^{\mathrm{2pt}}$, which are then used in the spectral decomposition of $r$ to fit the matrix elements. For SumFit, we use $r$ at $t_{s}=4a,6a,8a,10a$, summing over $\tau\in[2a,t_{s}-2a]$, and extract the matrix element from the slope of a linear fit in $t_{s}$. All fits are performed with an uncorrelated $\chi^{2}$, with correlations preserved through configuration-wise bootstrap. More details can be found in Appendix A.
@@
Figure 1 shows representative results at $P_{z}=1.62~\mathrm{GeV}$, $b_{T}=8a$, and $b_{z}=0$, for the unpolarized and helicity matrix elements. The consistency across different methods indicates that excited-state effects are under control. For the final analysis, we average over all three methods on a bootstrap sample-by-sample basis. Additional details and examples are provided in Appendix A.
Figure 2: The $x$-dependent quasi-TMD beam functions are presented as functions of $x$ at several values of $b_{T}$. From left to right, the panels the isoscalar unpolarized, isovector unpolarized, and isovector helicity cases, respectively.